OpenStudy (anonymous):

pls. help/teach me find the 2nd & 3rd derivative of this: A.) y=(4-x^2)^1/2 .....y=the square root of 4-x^2 B.) y=(x+5)^2

OpenStudy (anonymous):

\[A.) y=\sqrt{4-x ^{2}} B.) y=(x+5)^{2}\]

OpenStudy (loser66):

find second and third derivative of them? and they are separated problems, right?

OpenStudy (anonymous):

yea

OpenStudy (loser66):

do B) first, ok? it's easier. XD

OpenStudy (anonymous):

how

OpenStudy (loser66):

what do you want to go? combine or simplest form? I choose simplest. Ok, expand (x+5)^2=?

OpenStudy (loser66):

knock knock, do it, I guide you step by step. Cannot jump friend.

OpenStudy (anonymous):

how to expand it

OpenStudy (loser66):

how to expand (a+b)^2?

OpenStudy (anonymous):

so i'll use the foil method

OpenStudy (loser66):

it 's a perfect square!! It has formula to expand, friend. You study derivative , moreover you want to take second and third derivative but not know how to expand a perfect square? are you kidding me?

OpenStudy (loser66):

read and apply http://mathsfirst.massey.ac.nz/Algebra/SimpExp/CS.htm

OpenStudy (anonymous):

aˆ2 + 2ab + bˆ2

OpenStudy (loser66):

yes, but your is (x+5) ^2 , so??

OpenStudy (anonymous):

sorry i forgot

OpenStudy (anonymous):

x^2+25

OpenStudy (anonymous):

tnx for the link

OpenStudy (loser66):

where is the middle term?

OpenStudy (loser66):